# Overcoming the space-charge dilemma in low-energy heavy ion beams via a multistage acceleration lens system

### v1

@@ II The Space-Charge Dilemma in Heavy Ion Transport
Crucially, Eq. (1) also reveals a counter-intuitive “brightness paradox”: attempting to improve beam quality by reducing the initial radius (e.g., from 10 mm to 7 mm) actually amplifies the divergence (). Thus, simple geometric collimation is futile. To overcome these physical limitations, we propose the “Active Lensing” scheme [Fig. 1(c)]. By optimizing the voltage cofiguration (), we impart a strong curvature to the acceleration field, generating an active focusing force that continuously counteracts the space-charge repulsion throughout the transport line.
@@ III Design window of low energy beam transport (LEBT) for heavy ion beam
First, we estimate the beam behavior inside the multi-stage acceleration tube. We assume the injection of a gold negative ion beam from the source with an energy of 21 keV, 0 rad divergence, and a current of 100 A. Within an individual electrode gap ( m), the beam expansion is approximately mm, rendering the loss per gap negligible. However, when considering the total length of the multi-stage acceleration tube ( m), where the voltage transitions from 21 kV () to 64 kV (), the cumulative expansion reaches mm. This suggests that in the high-current regime, space-charge expansion cannot be ignored even within the acceleration tube, necessitating careful electrode design. In contrast, for hydrogen, the beam expansion remains negligible under the same conditions. However, it should be noted that when the beam current reaches the mA order, space-charge forces comparable to those of the gold beam become significant, necessitating similar caution.
@@ V Transmission efficiency through a tandem accelerator for Heavy Ion Beam
Figure 5 shows the experimental results of the beam transmission efficiency ( input-output ratio) through the tandem accelerator as a function of the injected beam current. We compared the performance of the conventional uniform acceleration (blue circles) with the optimized acceleration using Active Lensing (red circles) applied at the multi-stage acceleration electrodes prior to injection.

### v2

@@ II The Space-Charge Dilemma in Heavy Ion Transport
Crucially, Eq. (1) also reveals a counter-intuitive “brightness paradox”: attempting to improve beam quality by reducing the initial radius (e.g., from 10 mm to 7 mm) actually amplifies the divergence (). Thus, simple geometric collimation is futile. To overcome these physical limitations, we propose the “Active Lensing” scheme [Fig. 1(c)]. By optimizing the voltages and to their optimal values and , a strong curvature is imparted to the accelerating field, generating an active focusing force that continuously counteracts space-charge repulsion throughout the transport line.
@@ III Design window of low energy beam transport (LEBT) for heavy ion beam
First, we estimate the beam behavior inside the multistage acceleration tube. We assume the injection of a gold negative ion beam from the source with an energy of 21 keV, 0 rad divergence, and a current of 100 A. Within an individual electrode gap ( m), the beam expansion is approximately mm, rendering the loss in the gap negligible. However, when considering the total length of the multistage acceleration tube ( m), during which the acceleration voltage gradually increases from 21 kV to 64 kV, the cumulative expansion reaches mm. This suggests that in the high-current regime, space-charge expansion cannot be ignored even within the acceleration tube, necessitating careful electrode design. In contrast, for hydrogen, the beam expansion remains negligible under the same conditions. However, it should be noted that when the beam current reaches the mA order, space-charge forces comparable to those of the gold beam become significant, necessitating similar caution.
@@ V Transmission efficiency through a tandem accelerator for Heavy Ion Beam
Figure 5 shows the experimental results of the beam transmission efficiency ( input-output ratio) through the tandem accelerator as a function of the injected beam current. We compared the performance of the conventional uniform acceleration (blue circles) with the optimized acceleration using Active Lensing (red circles) applied at the multistage acceleration electrodes prior to injection.